Demand for earbuds is represented by: P = 402 – 2Q The costs are $8 per set of earbuds. In order to stock earbuds, other items which average $2 in profit must be taken off the shelf.
How many earbuds should you order?
Answer #1

MC= 8+2 =10

MR is twice the slope of the demand curve

MR = 402-4Q

Setting MC=MR for profit maximization

402-4Q=10

402-10 = 4Q

Q = 392/4 = 98